OpenStudy (anonymous):

FREE MEDAL!! Let f(x) = x2 - 16. Find f-1(x).

OpenStudy (anonymous):

would it be the sqrt of x+16????

OpenStudy (anonymous):

\[f(x) = 2x - 16\] \[f^-1(x) = \frac{ 1 }{ 2x-16 }\] \[= \frac{ 1 }{ 2(x-8) }\]

OpenStudy (anonymous):

or \[=\frac{ 1 }{ 2 }(x-8)^-1\] depending on how your teacher/book prefers the answer be displayed

OpenStudy (anonymous):

4square root of x square root of x plus 16 x squared over 16 1 over quantity of x squared minus 16 those are the answer choices